OpenStudy (anonymous):

How to isolate the x and y variables in this equation: (x+5)/((x-3)(x+1))

myininaya (myininaya):

i dont see y

OpenStudy (anonymous):

Oh I forgot, it is suppose to be y=(x+5)/((x-3)(x+1)).

OpenStudy (anonymous):

this sure looks just like the "even odd" problem i woke up to. is it?

OpenStudy (anonymous):

surely you have not been here all this time, have you?

myininaya (myininaya):

me no? i was helping someone else i didn't know he had replied back lol

OpenStudy (anonymous):

Yeah, I'm suppose to find the horizontal asymptote by solving for x in terms of y. No I just logged in.

OpenStudy (anonymous):

By the way, thanks for your help in the other question. It really helped me see the difference between odd and even functions.

myininaya (myininaya):

hey you don't need to solve for x to find horiztonal asymptote

myininaya (myininaya):

all you need to do is follow this: if degree of top=degree of bottom, then take the coefficient of the term with highest exponent in top and divide by the coefficient of the term with the highest exponent in the bottom. for example f(x)=(x^2-1)/(2x^2-3), horizontal asymptote is y=1/2 if degree of top>degree of bottom, then you do not have a horizontal aymptote example: f(x)=(x^3-4)/(x-1) then there is no horiztonal asymptote if degree of top<degree of bottom, then the horizontal asymptote is y=0 example: f(x)=(x+5)/((x-3)(x+1))